# Zero-shot LLM-guided Counterfactual Generation for Text

### v1

@@ 5 Counterfactual Explanations via LLM-generated Counterfactuals
instance to observe a diverse outcome." Enclose the generated text within <new> tags.
@@
Table 1: Prompt structure for the two variants of our pipeline when used for counterfactual explanations. <task on task-dataset> denotes the task description such as “natural language inference on the SNLI dataset". <$x_{i}$> denotes the input text. <$y_{i}$> denotes the label output by the black-box classifier for the input text <$x_{i}$>. <$y_{cf}$> denotes the desired label for the counterfactual text (any label other than <$y_{i}$>).
5.1 Methodology
To generate counterfactual explanations for a black-box text classifier $f(\cdot)$ that predicts $f(x_{i})=y_{i}$, we use the tuple $(x_{i},y_{i})$ in the generation step, thereby replacing the ground truth label in Figure 2 by the model-predicted label, since we aim to explain why the model predicted $y_{i}$ for the input sample $x_{i}$. In our experiments, we use a DistilBERT model  [38] , fine-tuned on the specific task dataset as the black-box model we aim to explain. Note that since our counterfactual generation process is model-agnostic, the same procedure can be applied to any black-box classifier in place of DistilBERT. Inspired by prior work [5], we develop and experiment with two variants of FIZLE: (1) FIZLEnaive: which directly generates the counterfactual explanation, and (2) FIZLEguided: which first extracts words that may have caused the predicted label, and then uses those selected words to generate a counterfactual explanation, in a two-step manner. We hypothesize that the two-step generation may result in more effective and better quality counterfactual explanations, due to the additional guidance provided to the LLM, analogous to prior work such as Chain of Thought [47]. We show the prompts used in both the variants in Table 1.
5.2 Evaluation Metrics
@@
We use Label Flip Score to measure the effectiveness of the generated counterfactual explanations. For each input text $x_{i}$ in the test split of the dataset, with correctly predicted label $f(x_{i})=y_{k}$, we evaluate the corresponding LLM-generated counterfactual $x_{i}^{cf}$ using the same black-box classifier $f(\cdot)$ and obtain a label for the counterfactual. For an effective counterfactual, the obtained label should be different from the original label $y_{k}$. Then Label Flip Score (LFS) is computed as:
@@
5.3 Baselines
Similar to other counterfactual generation methods [26, 48], we compare our proposed FIZLE pipeline with three representative baselines from three categories of similar works: (i) BAE [17] is a recent adversarial attack method that uses masked language modeling with BERT to perturb the input text by replacing masked words; (ii) CheckList  [36] is a method for behavioral testing of NLP models via test cases generated by template-based methods as well as masked language models like RoBERTa; (iii) Polyjuice  [48] is a recent counterfactual generation method that uses an auxiliary language model (such as GPT-2) to generate diverse counterfactuals. Note that unlike these baselines, our FIZLE pipeline does not require any auxiliary model, language model or dataset, thereby enabling a completely zero-shot generation.
5.4 Results: Effectiveness of Generated Counterfactual Explanations
@@
text-davinci-003
$(guided)$
95.78 0.87 0.186 54.9 0.84 0.153 39.6 0.83 0.179
text-davinci-003
$(naive)$
95.39 0.88 0.219 66.4 0.88 0.252 40.2 0.89 0.218
ChatGPT-3.5
@@
99.6 0.87 0.226 84.39 0.65 0.278 78 0.88 0.152
@@
Table 2: Evaluation results of both variants of our FIZLE framework in comparison to baselines: BAE [17], CheckList [36] and Polyjuice [48]. We report the Label Flip Score (LFS), semantic similarity (Sem. Sim) and normalized Levenshtein distance (Edit Dist.). Best LFS scores for each dataset are underlined.
@@
Following the experimental setup described above, we evaluate the generated counterfactuals to explain black-box classifiers for the three datasets, and compare to baselines. We show these quantitative results in Table 2. For each LLM, we evaluate both variants of our framework: FIZLEguided and FIZLEnaive. For effective and good quality counterfactual explanations, ideally we would expect high values of LFS and semantic similarity with low values of edit distance. Overall, we see varied performance of the LLMs and the two variants across the different tasks. Similar to other counterfactual generation works [26], we see an obvious trade-off between the Label Flip Score and the semantic similarity. This is intuitive since the more the generated counterfactual deviates from the original input text, higher the chances are for it to be a successful counterfactual for the original input (i.e, it would result in a label flip). Among the three baselines, we see CheckList fails completely in generating counterfactual explanations. We see satisfactory performance by BAE, except for the AG News dataset. For Polyjuice, even though the LFS scores are high, the poor textual similarity scores imply that the counterfactuals generated are not good quality and deviate from the input text significantly.
@@
For most of the LLMs we evaluated, FIZLEguided outperforms FIZLEnaive on the IMDB and AG News datasets. This may imply that the additional ‘guidance’ provided by identifying the input words before the counterfactual explanation generation step enables the generation of better counterfactuals. Interestingly, we do not see this trend for GPT-4, where the $naive$ variant performs better than the $guided$ one. This might be due to the fact that GPT-4 is extremely good at understanding complex tasks and instructions [8], and the additional feature extraction step in the $guided$ variant does not provide additional useful guidance, and perhaps even confuses the LLM. GPT-4 when used in the $naive$ variant of our pipeline, has the best performance for zero-shot counterfactual explanation generation, in terms of LFS. For natural language inference on the SNLI dataset, we see all LLMs struggle to generate good counterfactual explanations. GPT-4 performs well, possibly owing to its instruction and textual understanding capabilities [8], but the best performance is by the Polyjuice baseline. This poor performance of LLMs particularly on the SNLI dataset is further evidence towards LLMs struggling with inference and reasoning. This gap in the capabilities of recent LLMs on reasoning tasks has been observed by several recent efforts as well [34, 43].
@@
Lastly, we see the open-source models Llama 2 7B and 13B struggle to generate zero-shot counterfactual explanations with small number of edits, thus resulting in very high edit distances. The Llama 2 models struggle to keep the generated counterfactuals semantically similar to the original input, implying they either make too many edits to the input text, or output some unrelated, low-quality text that does not conform to the instructions provided in the prompt.
@@
Table 3: Performance of FIZLE-generated counterfactuals as contrast sets. C.s. Acc. refers to accuracy on the generated contrast sets, Original Test Acc. is the test accuracy on the corresponding paired original samples. Sem. Sim. refers to semantic similarity as computed by Eq. 2, Edit. Dist. is the token level distance as computed by Eq. 3, Cons. (%) is the consistency as computed by Eq. 4.
6 Evaluating Models via LLM-generated Counterfactuals
@@
Deep learning models such text classification models are often trained in a supervised manner using labeled training sets, and then evaluated on a hold-out test set. Such train-test splits of data usually arise from the same corpus that has same or similar sources and annotation guidelines. Therefore, in essence, standard evaluation using such hold-out test sets measure merely the in-distribution performance of the model, while in reality, the same model may demonstrate sub-par performance on out-of-distribution or in-the-wild test data [16]. To alleviate this issue to some degree, approaches such as evaluating using challenge sets or robustness to label-preserving perturbations, etc. have been explored by the community. One specific method of stress-testing such models is via contrast sets  [16]. A contrast set $C(x)$ is essentially a sample of points around a data point $x$, that is close to the local ground truth decision boundary. Samples in $C(x)$ may have same or different ground truth label as $x$. In practice, $C(x)$ can be a set of samples that are ‘close’ to $x$, i.e., have minimal edit distance from $x$, yet be ‘challenging’ for a trained model to classify. In the original contrast sets work, the authors advocate for an evaluation paradigm where dataset authors themselves create and release such contrast sets for model evaluation. However, we note that this is highly infeasible in practice, given the cost of expert creation of such challenging data points. Therefore, automated methods for designing such challenging evaluation sets in the form of contrast sets are highly desirable, albeit at the expense of trading off expert insights. One such automated method for developing contrast sets to evaluate models is that of counterfactual examples, as demonstrated by previous work [48]. Motivated by the effectiveness of counterfactuals as contrast sets in prior work [48], we envision the use of LLM-generated contrast sets as well for the same purpose of model evaluation. Here we describe the methodology for the generation and evaluation of such contrast sets using LLMs in a zero-shot manner.
6.1 Methodology
@@
For generating the contrast sets, we use the same LLMs as used in Section 5, except text-davinci-003111111since OpenAI deprecated this model by the time we performed these contrast set generation experiments., and prompt the LLM to generate counterfactuals in a zero-shot manner using the input text and ground truth label tuple $(x_{i},\hat{y}_{i})$. Unlike  [48], we do not use human annotators to label the generated counterfactuals. Therefore, differing from  [48], we only focus on counterfactuals that have the same label as the original input, and use these as contrast sets. We make this choice since the lack of human annotation and lack of step-by-step guidance (such as in FIZLEguided) would make it harder to validate whether the edits performed by the LLM are actually label flipping or not. Instead, we guide the generation process via the instruction in the prompt. We use the following prompt to perform the generation:
@@
6.2 Evaluation Metrics
For evaluating the goodness of the generated counterfactuals as contrast sets, we compare the accuracy of the target model $f(\cdot)$ on both the original test set and the generated contrast set. We also measure the consistency,
@@ 6.2 Evaluation Metrics
6.3 Baselines
6.4 Results: Effectiveness of Counterfactuals as Contrast Sets
We use the same DistilBERT models as in Section 5 that are fine-tuned for each of the 3 tasks (IMDB, SNLI and AG News). We evaluate each of these 3 models on both the original test set and the counterfactual one (i.e., the contrast sets) and show these results in Table 3. We obtain the performance values for Polyjuice-generated contrast sets from the original paper  [48]. For the ‘Expert’ baseline, the IMDB contrast sets are created by human experts in  [16]. Unfortunately, there does not exist any expert created contrast set for SNLI and AG News datasets. As evident from the test accuracies on both the original test set and the counterfactual one, we see a consistent decrease in performance on the generated counterfactuals over the original samples. The performance drop for the AG News dataset seems to be the least while interestingly, we see the highest performance drop on contrast sets for the SNLI dataset. Furthermore, we see GPT-3.5 and GPT-4 are able to create contrast sets with high degree of semantic similarity and low edit distance, thus being more desirable over Llama 2 generated contrast sets. Overall, the drop in accuracy and the consistency values seem analogous to similar results in literature (average drop in classification accuracy of around 6.8% according to  [48]) that use human-generated contrast sets for evaluation  [16, 48].
While this is promising, we do note the ethical concerns surrounding this: LLM-generated contrast sets may induce pre-existing biases that can propagate further bias and errors through evaluation and subsequent model improvement steps. One way to effectively use LLM-generated contrast sets is by broadly identifying the failure models of the model via probing the model using the LLM-generated contrast sets, and then employing human annotators or data creators to hone in on that specific failure mode to either generate more contrast sets or counterfactually augmented training data to fill the identified gap. Such a combined method would greatly reduce costs while still being effective in terms of model evaluation and development.
7 Conclusion & Future Work
@@
In this paper, we tackle the novel task of zero-shot general purpose counterfactual generation. To this end, we propose the paradigm of leveraging large language models as pseudo-oracles in order to generate plausible, high quality, and effective counterfactuals. Such a framework is zero-shot, requiring no training data or auxiliary models or even in-context learning. We validate our hypothesis via demonstrating the use and effectiveness of LLM-generated counterfactuals on two broad NLP model development tasks: (1) counterfactual explanation of black-box text classifiers, and (2) evaluation of black-box text classification models via contrast sets. Our results are promising and we see benefits to using our proposed FIZLE framework across the two use-cases and three downstream tasks. Our findings suggest the effective use of LLM-generated counterfactuals for explaining black-box NLP models, as well as potentially identifying failure models of NLP models via evaluation with contrast sets. We further discuss implications and how hybrid human-and-AI methods may benefit from our exploration.
@@
Future work may investigate modifications to this proposed framework such as generating contrast sets with opposite labels. More effort can also be put into validating the faithfulness of the generated counterfactual explanation and correctness of generated contrast sets. Another gap that can be addressed is how to improve the reasoning capabilities of the LLMs used in the pipeline, to improve explanation performance on the NLI task. Since one of the challenges in our method was to ensure that the generated text is actually a counterfactual, devising ways and human annotations to ensure more conformity of the generated counterfactual explanations to the definition of ‘counterfactual explanation’ is also be an area that can be improved. Finally, apart from these two use-cases of counterfactuals, LLM-generated counterfactuals can also be evaluated in tasks such as model training or improvement, uncovering biases in model predictions, incorporating fairness into model predictions, etc.
{ack}

### v2

@@ V Counterfactual Explanations via LLM-generated Counterfactuals
instance to observe a diverse outcome.” Enclose the generated text within <new> tags.
@@
V-A Methodology
To generate counterfactual explanations for a black-box text classifier $f(\cdot)$ that predicts $f(x_{i})=y_{i}$, we use the tuple $(x_{i},y_{i})$ in the generation step, thereby replacing the ground truth label in Figure 2 by the model-predicted label, since we aim to explain why the model predicted $y_{i}$ for the input sample $x_{i}$. In our experiments, we use a DistilBERT model  [46] , fine-tuned on the specific task dataset as the black-box model we aim to explain. Note that since our counterfactual generation process is model-agnostic, the same procedure can be applied to any black-box classifier in place of DistilBERT. Inspired by prior work [7], we develop and experiment with two variants of FIZLE: (1) FIZLEnaive: which directly generates the counterfactual explanation, and (2) FIZLEguided: which first extracts words that may have been responsible for the predicted label, and then uses those selected words to generate a counterfactual explanation, in a two-step manner. We hypothesize that the two-step generation may result in more effective and better quality counterfactual explanations, due to the additional guidance provided to the LLM, analogous to prior work such as Chain of Thought prompting [47]. We show the prompts used in both the variants in Table V.
V-B Evaluation Metrics
@@
[…]
@@
V-C Baselines
Similar to other counterfactual generation methods [6, 5], we compare our proposed FIZLE pipeline with three representative baselines from three categories of similar works: (i) BAE [50] is a recent adversarial attack method that uses masked language modeling with BERT to perturb the input text by replacing masked words; (ii) CheckList  [51] is a method for behavioral testing of NLP models via test cases generated by template-based methods as well as masked language models like RoBERTa; (iii) Polyjuice  [5] is a recent counterfactual generation method that uses an auxiliary language model (such as GPT-2) to generate diverse counterfactuals. Note that unlike these baselines, our FIZLE pipeline does not require any additional dataset or training, thereby enabling a completely zero-shot generation.
V-D Results: Effectiveness of Generated Counterfactual Explanations
TABLE I: Evaluation results of both variants of our FIZLE framework in comparison to baselines: BAE [50], CheckList [51] and Polyjuice [5]. We report the Label Flip Score (LFS), semantic similarity (Sem. Sim) and normalized Levenshtein distance (Edit Dist.). Best LFS scores for each dataset are in bold, second best is underlined.
@@ V-D Results: Effectiveness of Generated Counterfactual Explanations
GPT-3.5
@@
99.6 0.87 0.226 84.39 0.65 0.278 78.0 0.88 0.152
GPT-4o
$(guided)$
93.56 0.85 0.389 62.5 0.74 0.1917 66.56 0.92 0.092
GPT-4o
$(naive)$
99.57 0.72 0.56 90.12 0.41 0.549 79.39 0.90 0.157
GPT-4o-mini
$(guided)$
98.58 0.82 0.413 66.33 0.76 0.173 47.09 0.81 0.185
GPT-4o-mini
$(naive)$
100.0 0.63 0.642 66.6 0.59 0.485 51.4 0.83 0.282
@@
Llama 3 8B
$(guided)$
90.95 0.80 0.453 47.48 0.73 0.197 44.89 0.81 0.216
Llama 3 8B
$(naive)$
97.34 0.71 0.522 75.36 0.71 0.505 61.6 0.71 0.466
@@
Following the experimental setup described above, we evaluate whether generated counterfactuals can be used to explain black-box classifiers for the three datasets, as compared to the baselines. We show these quantitative results in Table I. For each LLM, we evaluate both variants of our framework: FIZLEguided and FIZLEnaive. For effective and good quality counterfactual explanations, ideally we would expect high values of LFS and semantic similarity with low values of edit distance. Overall, we see varied performance of the LLMs and the two variants across the different tasks. Similar to other counterfactual generation works [6], we see an obvious trade-off between the Label Flip Score and the semantic similarity. This is intuitive since the more the generated counterfactual deviates from the original input text, higher the chances are for it to be a successful counterfactual for the original input (i.e, it would result in a label flip). Among the three baselines, we see CheckList fails completely in generating counterfactual explanations. We see satisfactory performance by BAE, except for the AG News dataset. For Polyjuice, even though the LFS scores are high, the unsatisfactory textual similarity scores imply that the counterfactuals generated are not good quality and deviate from the input text significantly.
@@
Overall we do not see a clear winner between the two variants: FIZLEguided outperforms FIZLEnaive. However, while the LFS scores are better in the naive variant over the guided, the LLM is unable to preserve the textual similarity in comparison to the guided case. This may imply that the additional ‘guidance’ provided by identifying the input words before the counterfactual explanation generation step enables the generation of counterfactuals closer in semantics to the original input. GPT-4o and GPT-4o-mini when used in the $naive$ variant of our pipeline, have the best performance for zero-shot counterfactual explanation generation, in terms of LFS, for IMDB and AG News datasets, respectively. For natural language inference on the SNLI dataset, we see all LLMs struggle to generate good counterfactual explanations. GPT-4o performs well, possibly owing to its instruction and textual understanding capabilities [10], but the best performance is by the Polyjuice baseline. This poor performance of LLMs particularly on the SNLI dataset is further evidence towards LLMs struggling with inference and reasoning. This gap in the capabilities of recent LLMs on reasoning tasks has been observed by several recent efforts as well [52, 53].
@@
Lastly, we see the open-source models Llama 2 7B and 13B struggle to generate zero-shot counterfactual explanations with small number of edits, thus resulting in very high edit distances. The Llama 2 models struggle to keep the generated counterfactuals semantically similar to the original input, implying they either make too many edits to the input text, or output some unrelated, low-quality text that does not conform to the instructions provided in the prompt. However, the newer Llama 3 8b model outperforms both these models in most settings.
TABLE II: Performance of FIZLE-generated counterfactuals as contrast sets. C.s. Acc. refers to accuracy on the generated contrast sets, Original Test Acc. is the test accuracy on the corresponding paired original samples. Sem. Sim. refers to semantic similarity as computed by Eq. 2, Edit. Dist. is the token level distance as computed by Eq. 3, Cons. (%) is the consistency as computed by Eq. 4.
@@
Llama 3 8B 89.78 0.267 0.861 88.03 72.75 0.158 0.923 67.51 93.64 0.229 0.874 91.95
VI Evaluating Models via LLM-generated Counterfactuals
@@
Deep learning models such as text classification models are often trained in a supervised manner using labeled training sets, and then evaluated on a hold-out test set. Such train-test splits of data usually arise from the same corpus that has the same or similar sources and annotation guidelines. Therefore, in essence, standard evaluation using such hold-out test sets measure merely the in-distribution performance of the model, while in reality, the same model may demonstrate sub-par performance on out-of-distribution or in-the-wild test data [8]. To alleviate this issue to some degree, approaches such as evaluating using challenge sets or robustness to label-preserving perturbations, etc. have been explored by the community. One specific method of stress-testing such models is via contrast sets  [8]. A contrast set $C(x)$ is essentially a sample of points around a data point $x$, that is close to the local ground truth decision boundary. Samples in $C(x)$ may have same or different ground truth label as $x$. In practice, $C(x)$ can be a set of samples that are ‘close’ to $x$, i.e., have minimal edit distance from $x$, yet be ‘challenging’ for a trained model to classify. In the original contrast sets work [8], the authors advocate for an evaluation paradigm where dataset authors themselves create and release such contrast sets for model evaluation. However, we note that this is highly infeasible in practice, given the cost of expert creation of such challenging data points. Therefore, automated methods for designing such challenging evaluation sets in the form of contrast sets are highly desirable, albeit at the expense of trading off expert insights. One such automated method for developing contrast sets to evaluate models is that of counterfactual examples, as demonstrated by previous work [5]. Motivated by the effectiveness of counterfactuals as contrast sets in prior work [5], we envision the use of LLM-generated contrast sets as well for the same purpose of model evaluation. Here we describe the methodology for the generation and evaluation of such contrast sets using LLMs in a zero-shot manner.
VI-A Methodology
@@
For generating the contrast sets, we use the same LLMs as used in Section V, and prompt each LLM to generate counterfactuals in a zero-shot manner using the input text and ground truth label tuple $(x_{i},\hat{y}_{i})$. Unlike  [5], we do not use human annotators to label the generated counterfactuals. Therefore, differing from  [5], in our evaluation, we only focus on counterfactuals that have the same label as the original input, and use these as contrast sets. We make this choice since the lack of human annotation and lack of step-by-step guidance (such as in FIZLEguided) would make it harder to validate whether the edits performed by the LLM are actually label flipping or not. Instead, we guide the generation process via the instruction in the prompt. We use the following prompt to perform the generation:
@@
VI-B Evaluation Metrics
For evaluating the goodness of the generated counterfactuals as contrast sets, we compare the accuracy of the target model $f(\cdot)$ on both the original test set and the generated contrast set. Following  [5], we also measure the consistency,
$consistency=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[f(x_{i})=\hat{y}_{i}\wedge f(x_{i}^{cs})=\hat{y}^{cs}_{i}]\times 100$ (4)
VI-C Baselines
VI-D Results: Effectiveness of Counterfactuals as Contrast Sets
We use the same DistilBERT models as in Section V that are fine-tuned for each of the 3 tasks (IMDB, SNLI and AG News). We evaluate each of these 3 models on both the original test set and the counterfactual one (i.e., the contrast sets) and show these results in Table II. We obtain the performance values for Polyjuice-generated contrast sets from the original paper  [5]. For the ‘Expert’ baseline, the IMDB contrast sets are created by human experts in  [8]. Unfortunately, there does not exist any expert created contrast set for SNLI and AG News datasets. As evident from the test accuracies on both the original test set and the counterfactual one, we see a consistent decrease in performance on the generated counterfactuals over the original samples. The performance drop for the AG News dataset seems to be the least while interestingly, we see the highest performance drop on contrast sets for the SNLI dataset. Furthermore, we see GPT-3.5 and GPT-4 are able to create contrast sets with high degree of semantic similarity and low edit distance, thus being more desirable over Llama 2 generated contrast sets. Interestingly, we see that GPT-4o and GPT-4o-mini contrast sets often have better performance than the original test set, which could imply that these contrast sets are too ‘easy’, and therefore not functional. Overall, for the other LLMs, the drop in accuracy and the consistency values seem analogous to similar results in literature (average drop in classification accuracy of around 6.8% according to  [5]) that use human-generated contrast sets for evaluation  [8, 5]. Evaluating models with such LLM-generated contrast sets may thereby allow the model developer to investigate what type of samples the model is failing on, thereby informing choices regarding further robustness training.
While this is promising, we do note the ethical concerns surrounding this: LLM-generated contrast sets may induce pre-existing biases that can propagate further bias and errors through evaluation and subsequent model improvement steps. One hybrid way to effectively use LLM-generated contrast sets is by broadly identifying the failure models of the model via probing the model using the LLM-generated contrast sets, and then employing human annotators or data creators to hone in on that specific failure mode to either generate more contrast sets or counterfactually augmented training data to fill the identified gap. Such a combined method would greatly reduce costs while still being effective in terms of model evaluation and development.
VII Case Study: How has the performance of LLMs evolved over time?
In order to use LLMs for zero-shot counterfactual generation for stress-testing models, one needs to account for the fast evolving landscape of LLM development and release. These models, especially proprietary ones, undergo frequent model updates. Older models are often deprecated, making adoption of such solutions challenging without thorough re-evaluation of pipelines such as ours. To explore the effect of model release and deprecation life cycles, we conduct a small case study to evaluate how performance of this method of generating counterfactuals vary over time. We do this by reporting naive generation results for counterfactual explanations for GPT family of models. Models we use here, from oldest (2022) to newest (2024), are: text-davinci-003151515https://platform.openai.com/docs/deprecations $\rightarrow$ GPT-3.5 $\rightarrow$ GPT-4 $\rightarrow$ GPT-4o. We show this comparison in Figure 3. Keeping the expected trade-off between LFS and semantic similarity in mind, we see that for a task like NLI, newer models in the GPT family perform better, achieving high LFS scores while retaining semantic similarity, whereas for AG News, performance is varied: recent models like GPT-4 and GPT-4o achieve higher LFS but at the cost of textual similarity. While a more thorough evaluation is required in order to make broad claims, however, given these tasks and models, recent models like GPT-4o seem to be better at more complex tasks such as reasoning and inference, while a relatively older model like GPT-4 may be sufficient for other types of tasks.
@@ VII Case Study: How has the performance of LLMs evolved over time?
VIII Conclusion & Future Work
@@
In this paper, we explore the possibility of using LLM-generated zero-shot counterfactuals for stress-testing black-box text classification models. To this end, we propose a pipeline, with two variants, for generation of such counterfactuals in a zero-shot manner. We conduct experiments with a variety of proprietary and open LLMs, and evaluate generated counterfactuals on two broad NLP model development tasks: (1) counterfactual explanation of black-box text classifiers, and (2) evaluation of black-box text classification models via contrast sets. Our results are promising and we see benefits to using our proposed FIZLE pipeline across the two use-cases and three downstream tasks. Our findings suggest the effective use of LLM-generated counterfactuals for explaining black-box NLP models, as well as potentially identifying failure models of NLP models via evaluation with contrast sets. We further discuss implications and how hybrid human-and-AI methods may benefit from our exploration, along with looking at how model updates over time can affect the quality and performance of generated counterfactuals.
@@
Future work may investigate modifications to the generation pipeline such as plugging in an additional model, perhaps a small language mode (SLM) to evaluate label of generated contrast sets. More effort can also be put into validating the faithfulness of the generated counterfactual explanation and correctness of generated contrast sets. Further exploration could look into categorizing the failure modes of the black-box models, after being evaluated by these generated counterfactuals. This information can then be used to collect or generate more data for robustness training of these models. Since one of the challenges in our method was to ensure that the generated text is actually a counterfactual, devising ways and human annotations to ensure more conformity of the generated counterfactual explanations to the definition of ‘counterfactual explanation’ is also be an area that can be improved. Finally, apart from these two use-cases of counterfactuals, LLM-generated counterfactuals can also be evaluated in tasks such as model training or improvement, uncovering biases in model predictions, incorporating fairness into model predictions, etc.
Acknowledgments